you are given demand and supply schedule o burgers. where px is the price of a burger in rupees , QDX is the quantity demand and QSX is the quantity supplied.


px 100 200 300 400

QDX 1100 900 700 500

QSX 50 250 450 650


  1. CALCULATE DEMAND AND SUPPLY FUNCTION
1
Expert's answer
2021-08-12T08:55:52-0400

The quantity demanded of a good refers to the amount of the good the consumers of the good are willing and able to buy at some particular market price during a given period of time. The demand function of a good shows the quantity demanded of the good at various possible market prices. As we know when the price of the good is equal to 100 then the quantity demanded of the good is equal to 1100 so the first price quantity combination(P0, Q0) becomes equal to (100,1100) and when the price of the good increases and becomes equal to 200 then the quantity demanded of the good becomes equal to 900 so the second price quantity demanded combinations(P1, Q1) becomes equal to (200,900). So by using the above information, we calculate the demand function of the good as follows

"Q−Q0 =\frac{Q1−Q0}{P1−P0}(P−P0)\\Q-1100=\frac{ 900−1100}{200−100}(P−100)\\Q-1100= −2P +200\\Q= 1300−2P ..."1""


The equation "1" above shows the demand function of the good in the question.

The quantity supplied of a good refers to the amount of the good the sellers of the good are willing and able to supply in the market at some particular market price during a given period of time. The supply function of a good shows the quantity supplied of the good in the market at various possible market prices. As we know when the price of the good is equal to 100 then the quantity supplied of the good becomes equal to 50 so the first price quantity combination(P0, Q0) becomes equal to (100,50), and when the price of the good increases and becomes equal to 200 then the quantity supplied of the good becomes equal to 250 so the second price quantity supplied combinations(P1, Q1) becomes equal to (200,250). So by using the above information we calculate the supply function of the good equation as follows

"Q−Q0=\frac{Q1−Q0}{P1−P0}(P−P0)\\Q-50=\frac{250−50}{200−100}(P−100)\\Q-50= 2(P−100) \\Q=2P−150 ..."2""

The equation "2" shows the supply function of the good.